### v1

@@ 4.1.1 Technology
A representative firm operates a Cobb-Douglas technology and produces a perishable consumption good using labor and capital.
@@
As in the stochastic growth economy of Brock and Mirman, (1972), the firm faces competitive input and output markets. We choose consumption good as numéraire, hence price of consumption is normalized to unity. We denote the wage by $w_{t}$ and rental rate of capital by $r^{K}_{t}$.
@@ 4.1.3 Preferences
Households have time separable expected utility over consumption streams. We assume that the felicity function takes the constant relative risk aversion form with coefficient $\gamma$. Hence, household rank stochastic consumption stream according to:
@@ 4.1.4 Asset markets
Households are able to re-allocate their consumption across time and states by trading in two assets. Two assets and infinitely many possible shock realizations next period [^24] We assume innovations to the logarithm TFP and capital depreciation to be gaussian. imply incomplete markets environment. Furthermore, there are two additional financial frictions. First, households are subject to short sale limits on capital and bonds. Second, while the bond is a liquid asset, capital is an illiquid asset and households have to pay adjustment costs when adjusting their capital holdings.
The capital holdings adjustment costs are given by $\psi(k_{t+1}^{h+1},k_{t}^{h})=\xi^{\text{adj}}\left(k_{t+1}^{h+1}-k_{t}^{h}\right)^{2}$. Short sale constraints on both assets are specified as exogenous limits $b_{t+1}^{h+1}\geq\underline{b}$ and $k_{t+1}^{h+1}\geq\underline{k}$.
The net supply of bonds is fixed at $B$. Bonds can be purchased or sold at equilibrium price $p_{t}$, and they deliver a risk-free payoff of 1 in period $t+1$. Purchasing a unit of capital in period $t$, promises a risky payout of $1-\delta_{t+1}+r_{t+1}^{K}$ in period $t+1$. There are two sources of risk in returns to capital: first, the rental rate of capital $r_{t+1}^{K}$, depends on the total factor productivity $A_{t+1}$ in period $t+1$. Second, the depreciation rate of capital is stochastic as well and given by
@@
with $\mu^{\delta}<0$ and $\epsilon_{t+1}^{\delta}\sim\mathcal{N}(0,1)$. Normal and disaster times are modeled as two discrete Markov regimes with transition matrix $\Pi^{\text{regime}}=\begin{bmatrix}\pi^{n\rightarrow n}&1-\pi^{n\rightarrow n}\\ 1-\pi^{d\rightarrow d}&\pi^{d\rightarrow d}\end{bmatrix}$, where $\pi^{n\rightarrow n}$ and $\pi^{d\rightarrow d}$ denote the persistence of the normal and the disaster regime respectively.
@@ 4.1.5 Household problem
Recursively formulated, the households’ problem is characterized by the following Bellman equation.
@@
In every period, households receive labor and asset income, and decide how much to invest in each of the two assets, subject to adjustment costs and short-sale constraints. The time index and age superscript stand for the dependence on state variables.
The solution to the household problem is characterized by a set of Karush-Kuhn-Tucker (KKT) conditions for each age-group $h=1,\dots,H-1$ and each asset
@@
$0$ $\leq\lambda_{t}^{k,h}$ (37)
Making use of the Fischer-Burmeister function (see, Jiang,, 1999) $\psi^{FB}(a,b):=\sqrt{a^{2}+b^{2}}-a-b$, a first-order condition and the corresponding complementary slackness condition can be collapsed into a single equation for each age-group and each asset. Analogously to equation (11) for the representative agent model, we rearrange the equations, such that, in equilibrium, the equation has to evaluate to zero, and the deviations from zero are expressed in units of relative consumption error.
@@ 4.1.6 Equilibrium
is given by the regime indicator, $\chi_{t}$, total factor productivity $A_{t}$, capital depreciation $\delta_{t}$, and the distribution of assets across the age-groups $\mu_{t}=\{(b_{t}^{h},k_{t}^{h})\}_{h=1}^{H}$. [^25] Strictly speaking, because we assume that households enter the economy without assets, and because of market clearing, a sufficient statistic for the asset distribution consists of H1H-1 free parameters for capital and H2H-2 for the bond.
@@
The functional rational expectations equilibrium of our economy consists of $H$ policy functions $\pi^{k,h}(\mathbf{X}^{\text{state}}_{t})=k_{t+1}^{h+1}$ for investment in capital, $H$ policy functions $\pi^{b,h}(\mathbf{X}^{\text{state}}_{t})=b_{t+1}^{h+1}$ for investment in the bond, and a price function $\pi^{p}(\mathbf{X}^{\text{state}}_{t})=p_{t}$, consistent with the KKT conditions encoding household optimal behavior and a market clearing equation. [^26] We use the first-order conditions of representative firm directly to compute the equilibrium wage and rental rate as a function of the state of the economy. The aim of our algorithm is to compute approximations to these equilibrium functions.
@@ 4.2 Parameterization
We set the model parameters to standard values in the literature. One model period corresponds to one year, and we model $H=72$ overlapping generations. The model ages $h=1$ to $72$ corresponds to adult life from $21$ years to $92$ years. The age dependent efficient units of labor are chosen to generate a life-cycle profile of labor income and are shown in figure 3.
Figure 3: Efficiency units of labor over the lifecycle.

### v2

@@ 4.1.1 Technology
A representative firm operates a Cobb-Douglas technology and produces a perishable consumption good using labor and capital
@@
As in the stochastic growth economy of Brock and Mirman, (1972), the firm faces competitive input and output markets. We choose the consumption good as the numéraire, hence price of consumption is normalized to unity. We denote the wage by $w_{t}$ and rental rate of capital by $r^{K}_{t}$.
@@ 4.1.3 Preferences
Households have time separable expected utility over consumption streams. We assume that the felicity function takes the constant relative risk aversion form with coefficient $\gamma$. Hence, households rank stochastic consumption stream according to
@@ 4.1.4 Asset markets
Households are able to re-allocate their consumption across time and states by trading in two assets. Two assets and infinitely many possible shock realizations next period [^28] We assume innovations to the logarithm of TFP and capital depreciation to be Gaussian. imply an incomplete-markets environment. Furthermore, there are two additional financial frictions. First, households are subject to short sale limits on capital and bonds. Second, while the bond is a liquid asset, capital is an illiquid asset and households have to pay adjustment costs when adjusting their capital holdings. The adjustment costs are given by $\psi(k_{t+1}^{h+1},k_{t}^{h})=\xi^{\text{adj}}\left(k_{t+1}^{h+1}-k_{t}^{h}\right)^{2}$. Short sale constraints on both assets are specified as exogenous limits $b_{t+1}^{h+1}\geq\underline{b}$ and $k_{t+1}^{h+1}\geq\underline{k}$. The net supply of bonds is fixed at $B$. Bonds can be purchased or sold at equilibrium price $p_{t}$, and they deliver a risk-free payoff of 1 in period $t+1$. Purchasing a unit of capital in period $t$ promises a risky payout of $1-\delta_{t+1}+r_{t+1}^{K}$ in period $t+1$. There are two sources of risk in returns to capital: first, the rental rate of capital $r_{t+1}^{K}$ depends on the total factor productivity $A_{t+1}$ in period $t+1$. Second, the depreciation rate of capital in period $t+1$ is stochastic as well and given by
@@
with $\mu^{\delta}<0$ and $\epsilon_{t+1}^{\delta}\sim\mathcal{N}(0,1)$. Normal and disaster times are modeled as two discrete Markov regimes with transition matrix $\Pi^{\text{regime}}=\begin{bmatrix}\pi^{n\rightarrow n}&1-\pi^{n\rightarrow n}\\ 1-\pi^{d\rightarrow d}&\pi^{d\rightarrow d}\end{bmatrix}$, where $\pi^{n\rightarrow n}$ and $\pi^{d\rightarrow d}$ denote the persistence of the normal and disaster regimes, respectively.
@@ 4.1.5 Household problem
Recursively formulated, the households’ problem is characterized by the following Bellman equation
@@
In every period, households receive labor income and the payoff of their asset holdings, and decide how much to consume and how much to invest in each of the two assets, subject to adjustment costs and short-sale constraints. The time index denotes the dependence on the aggregate state variables.
The solution to the household problem is characterized by a set of Karush-Kuhn-Tucker (KKT) conditions. For each age-group $h=1,\dots,H-1$ and each asset, we have two sets of KKT conditions
@@
$0$ $\leq\lambda_{t}^{k,h}.$ (38)
For each age-group and each asset we use the Fischer-Burmeister function (see, Jiang,, 1999) $\psi^{FB}(a,b):=a-b-\sqrt{a^{2}+b^{2}}$ to collapse the set of KKT conditions into a single equation. The above KKT conditions are fully characterized by
$0$ $=\psi^{FB}\left(p_{t}u^{\prime}(c_{t}^{h})-\beta\text{E}\left[u^{\prime}(c_{t+1}^{h+1})\right],b_{t+1}^{h+1}-\underline{b}\right)$ (39)
$0$ $=\psi^{FB}\left(\left(1+\frac{\partial}{\partial k_{t+1}^{h+1}}\psi(k_{t+1}^{h+1},k_{t}^{h})\right)u^{\prime}(c_{t}^{h})\right$.
$\left.~~~-\beta\text{E}\left[u^{\prime}(c_{t+1}^{h+1})\left(1-\delta_{t+1}+r_{t+1}^{k}-\frac{\partial}{\partial k_{t+1}^{h+1}}\psi(k_{t+2}^{h+2},k_{t+1}^{h+1})\right)\right],(k_{t+1}^{h+1}-\underline{k})\right)$ (40)
Analogously to equation (12) for the representative agent model, we rearrange the equations so that the deviations from zero are expressed in units of relative consumption error.
@@ 4.1.6 Equilibrium
is given by the regime indicator, $\chi_{t}$, total factor productivity $A_{t}$, capital depreciation $\delta_{t}$, and the distribution of assets across the age groups $\mu_{t}=\{(b_{t}^{h},k_{t}^{h})\}_{h=1}^{H}$. [^29] Strictly speaking, because we assume that households enter the economy without assets, and because of market clearing, a sufficient statistic for the asset distribution consists of H1H-1 free parameters for capital and H2H-2 for the bond, because the bond supply is fixed.
@@
The functional rational expectations equilibrium of our economy consists of $H$ policy functions $\pi^{k,h}(\mathbf{X}^{\text{state}}_{t})=k_{t+1}^{h+1}$ for investment in capital, $H$ policy functions $\pi^{b,h}(\mathbf{X}^{\text{state}}_{t})=b_{t+1}^{h+1}$ for investment in the bond, and a price function $\pi^{p}(\mathbf{X}^{\text{state}}_{t})=p_{t}$, such that the KKT conditions hold and the bond and capital markets clear. [^30] We use the first-order conditions of representative firm directly to compute the equilibrium wage and rental rate as a function of the state of the economy. The goods market clears by Walras’ law. The aim of our algorithm is to compute approximations to these equilibrium functions.